P ( -2.81 < z < 0 ) 


1
Expert's answer
2022-03-18T05:28:40-0400

"P\left( -2.81<z<0 \right) =\varPhi \left( 0 \right) -\varPhi \left( -2.81 \right) =0.5-0.00248=0.49752"